I need help with these questions.

1. Calculate the force of graviational attraction between an SUV, mass-2000kg, and the Earth.

2. Calculate the gravitational field strength on the surface of Ganymede. How does this compare to the Earth? Our Moon?

3. Calculate the force of gravitational attraction between two carbon 16 atoms when theya re 1x10^-13 m apart. The mass of a proton or a neutron is 1.67x10^-27 kg, and the mass of an electron is 9.11x10^-31

4. Mary weighs 750N on Earth. She travels to the surface of Planet X, of radius=4.28x10^6 m. If the force of gravitational attraction between Mary and Planet X is 500N, calculate the mass of Planet X
Answer #1

1)

force of attraction = GMm/R*R = 6.67*10^-11*2000*5.97*10^24/(64*64*10^10) = 19443.3 N


2)The gravitational fiels strenth on ganymede surface will be given by F = GM/R^2

where G is gravitational constant , M is mass of ganymede and R is its radius.

The field strength is smaller than that of earth, because the mass of ganymede is much smaller than that of earth.

However its field is greater than that of moon because of the size.


3)

mass of 1 carbon 16 atom = (6+10)*1.67*10^-27 + 6*9.11*10^-31 = 2.67*10^-26 kg

So force = 6.67*10^-11*2.67*2.67*10^-52/(10^-26) = 4.75*10^-36 N


4)

force = GmM/R^2 = 6.67*10^-11*(750/9.8)*M/(4.28*4.28*10^12)


or M = 1.74 *10^24 Kg